How much energy (in calories) is needed to warm 75.0 grams of water from 14.4 degrees C to 19.2 degrees C?
1
Expert's answer
2017-04-28T14:28:08-0400
Q = cmΔT
ΔT = 19.2 – 14.4 = 4.8˚C = 4.8 K
C (water) = 4.218 kJ/(kg·K)
Q = 4.218 · 0.075 · 4.8 = 1.52 kJ